what is the next number in the series of 50,122,170,290,362,?

Grade:12

2 Answers

Ratnam Sarika
5 years ago
72+1=50    2.112+1=122.     3. 132+1=170.      4.172+1=290.  5.192+1=362. ….............6. 232+1=530                                                                                             
 Hence the next number in series is 530 .
HOPE     IT.     WOULD HELP YOU.
  1.  
dev gupta
5 years ago
the series is 7^2+ 1 , 11^2 +1 , 13^2+1, 17^2+1,....and so on
so we can see that 
now we will see the difference of the main number which are
7,11,13,17,19
now
11-7=4
13-11=2
17-13=4
19-17=2
and so on
so we cam to the result that the next number will be 23 in this sub series
so the next number of the series 
50,122,170,290,362 
will be 
23^2+1=529+1=530
50,122,170,290,362 .530
hope you will understand 
please reply if any problem.........best of luck